OpenStudy (anonymous):

how to convert the density of magnesium metal at 1.74 g/cm3 to kg/m3. I am a newbee! and any suggestions on conversion will help as well

OpenStudy (anonymous):

100 cm = 1 m , (100cm)^3 = (1m)^3 100g =1kg if you still need help ask. Hope this helped!

OpenStudy (kainui):

Think of conversion factors as multiplying times 1 in a new, abstract kind of way. Since we know 1000 grams is the exact same thing as 1 kilogram, you can think of 1kg/1000g as multiplying something by 1. It doesn't change the original quantity, it just changes the units since in division, the units will divide out instead of numbers. Here's an example:  See how the grams divide out, since grams over grams is like anything divided by itself, so it cancels out the same as 5/5 = 2/2 = 235/235 = karbledygook/karbledygook. Does that help make sense of conversion factors a little bit more or did I just confuse you more?